OpenStudy (christos):

Calculus 2 See below

OpenStudy (christos):

\[\begin{align} & if\int{\tan (x)dx=-\log (\cos (x))+C} \\ & thenalso\int{\tan (x)dx=-\ln (\cos (x))+C} \\ \end{align}\] True or false ?

hartnn (hartnn):

the log in the answer has base =e

OpenStudy (christos):

oh

OpenStudy (christos):

how did you know

OpenStudy (christos):

?

OpenStudy (christos):

Like wolfram doesn't show anything about it http://www.wolframalpha.com/input/?i=int+of+tan%28x%29

hartnn (hartnn):

integral f'(x) / f(x) dx = log f(x) +c here, it is assumed that log has base e

OpenStudy (christos):

oh ok thanks

hartnn (hartnn):

wolf says that log has natural base